If tan35=k then the value of tan145-tan125/1+tan145tan125=
 

Grade:11

1 Answers

Arun
6 years ago
Given that Tan 35 = k

⇒ Tan 145 - Tan 125 / 1 + Tan 145 * Tan 125

=
 Tan ( 145 - 125) = Tan 20

⇒ Tan 20 = Cot ( 90 - 20 ) = Cot 70

= Cot (2 x 35) = ( 1 - Tan35) / 2 Tan 35  = ( 1 - k2) / 2k.